# Geometry and cohomology of compactified Deligne–Lusztig varieties

### v2

@@ Abstract.
For connected reductive groups together with a Frobenius root $F$ that preserves the set of simple reflections, we show that for every tuple $(w_{1},\ldots,w_{r})$ of Weyl group elements there exists a minimal length element in an $F$-conjugacy class in $W$ such that the cohomology of the structure sheaf and the canonical sheaf for the associated compactified Deligne–Lusztig varieties are isomorphic respectively.
@@ 1.1. Background
Let $q$ be a power of a prime $p$ and let $k_{0}=\mathbb{F}_{q^{d}}$ for a positive integer $d$. Let $k$ be an algebraic closure of $\mathbb{F}_{p}$ containing $k_{0}$. Let $G_{0}$ be a connected reductive $k_{0}$-group with $q^{d}$-Frobenius $F_{0}:G_{0}\to G_{0}$. Let $G$ be the base change of $G_{0}$ to $k$ and let $F:G\to G$ be an endomorphism such that $F^{d}=F_{0}\otimes_{k_{0}}{\rm Id}$. The group of fixed points $G^{F}$ is known as a finite group of Lie type.
Let $T\subset B$ be an $F$-stable maximal torus and an $F$-stable Borel subgroup of $G$. Denote by $W:=N(T)/T$ the corresponding Weyl group and let $S\subseteq W$ be the set of simple reflections. For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$. Throughout the introduction, we fix a reduced expression of $w$ and denote by $\overline{X}(w)$ the corresponding smooth compactification of $X(w)$.
In [DL], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and certain Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero. The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic counterpart. However, for $G^{F}={\rm GL}_{2}(\mathbb{F}_{q})$ and $w$ the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
@@
In [Wang22], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of ${\rm GL}_{n}(\mathbb{F}_{q})$. Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S\ |\ s\leq w\}$. As a consequence of this theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lusztig variety $X(w)$ for ${\rm GL}_{n}(\mathbb{F}_{q})$.
@@ 1.2. Main results
The aim of this paper is to extend the main theorem of [Wang22] to all pairs $(G,F)$ for which $F(S)=S$. The assumption on $F$, for example, is satisfied when $G$ is split over $k_{0}$, or when $G^{F}$ is the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$.
If $G^{F}$ is a finite simple group of Lie type arising from a pair $(G,F)$ as in in Section 1.1, then the classification of such groups implies that $F(S)=S$ is satisfied. In fact, $F$ would be the Frobenius endomorphism (i.e. $d=1$) except for the cases for Suzuki and Ree groups, for which $d=2$ and additional assumptions on $q$ are required (Section 5.2).
In our setting, we encounter the minimal length elements in the $F$-conjugacy classes of $W$ (cf. Section 4.2), which is a special class of elements in $W$ with nice properties [GeckKimPfeifferMinlength, Theorem 2.6], [HeNieminlength, Theorem 3.1]. For example, if $G^{F}$ is ${\rm GL}_{n}(\mathbb{F}_{q})$ or the group of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$, then such elements are precisely the Coxeter elements in parabolic subgroups of $W$.
We show that the study of the cohomology of the structure sheaf and the canonical sheaf of $\overline{X}(w)$ for any $w\in W$ can always be reduced to that of a minimal length element in an $F$-conjugacy class in $W$.
Theorem 5.1.1.
Assume $F(S)=S$ and let $w\in W$. There exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
@@
Note that the $F$-conjugacy class $C$ in the theorem may not contain $w$. However, for any given $w$, one can determine $x_{0}$ and $C$ explicitly by following our recipe given in the proof of Theorem 5.1.1. We will also show that the cohomology groups in the theorem do not depend on the choice of reduced expressions of $w$ and $x_{0}$ and their corresponding smooth compactifications.
On the other hand, smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic. To avoid this ambiguity, we work with compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements, cf. Definition 2.1.1, [DMR, Definition 2.3.2]. This notion makes the construction of compactifications of Deligne–Lusztig varieties more straightforward and uniform. For example, setting $r=1$ yields the Zariski closure $X(\underline{w_{1}})=\overline{X(w_{1})}$ of $X(w_{1})$ in $G/B$. Setting $w_{1},\ldots,w_{r}\in S$ such that $w_{1}\cdots w_{r}\in W$ is a reduced expression yields the smooth compactification $X(\underline{w_{1}},\ldots,\underline{w_{r}})=\overline{X}(w_{1}\cdots w_{r})$ defined by Deligne and Lusztig.
In this generality, we give a connectedness criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [BonnafeRouquierOntheirred]. We also show that they have $F_{p}$-rational singularities (Section 2.4) and thus are normal, Cohen–Macaulay, and pseudo-rational. Furthermore, we provide a list of compactifications $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ of Deligne–Lusztig varieties which are smooth.
The structure of the paper is as follows. We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section 2. In Section 3, we show that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{i},i=1,\ldots,r,$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and the canonical sheaf on the associated compactified Deligne–Lusztig varieties.
Section 4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties. They are associated with $F$-cyclic conjugations in the Weyl group. In Proposition 4.1.4, we relate the cohomology of the structure sheaf (resp. canonical sheaf) between these compactified Deligne–Lusztig varieties.
Finally, we prove Theorem 5.1.1 and extend the result to arbitrary compactified Deligne–Lusztig varieties of the form $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in Section 5.
@@ 2. Compactified Deligne–Lusztig varieties
We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements. This construction has been treated in [DL, §9] and [DMR, §2]. We review here the definitions and properties that are relevant for our context. Moreover, by adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be connected and show that they are $F_{p}$-rational.
@@ 2.1. Definitions and basic properties
Let $G$, $F:G\to G$, $T\subset B$, $S\subset W$ be as in Section 1.1. In the literature, the endomorphism $F$ is referred to as a Frobenius root or a Steinberg endomorphism. Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
Recall that by the Bruhat decomposition, the diagonal $G$-action on the product $G/B\times G/B$ yields a bijection between the Weyl group $W$ and $G\backslash\left(G/B\times G/B\right)$. The orbit corresponding to $w\in W$ is denoted by ${O(w)}$. It can be regarded as a fibre bundle over the flag variety $G/B$ with fibres isomorphic to the Schubert cell $BwB/B$.
@@
Since $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$, cf. [Jantzen, F.23], it is an integral, separated, normal, Cohen–Macaulay $k$-scheme locally of finite type. A proof of normality can be found in [DMR, Lemma 2.2.13]. The Cohen–Macaulay property behaves well under base change [stacks-project, Lemma 045T], so $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is Cohen–Macaulay follows from the classical result that Schubert varieties are Cohen–Macaulay (see the introduction of [Hashimoto]).
Denote by $F:G/B\to G/B$ the map induced by the endomorphism $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph. Since $F^{d}$ is a Frobenius endomorphism, we know that $F$ is surjective, finite ([DigneMichelBook, Proposition 8.1.2]), separated ([GoerWedhAlgGeom1, Proposition 9.13]), and purely inseparable ([EGA1, (3.7.6)]). Moreover, the induced map $F:W\to W$ is a group automorphism [DigneMichelBook, Lemma 4.2.21].
@@
For any $w_{1},\ldots,w_{r}\in W$, the $k$-scheme $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$. If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$, cf. [DMR, Proposition 2.3.5, 2.3.6 (i)]. In general, compactified Deligne–Lusztig varieties have $F_{p}$-rational singularities, as we will see in Proposition 2.4.1.
@@
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=\bigcup_{w_{i}^{\prime}\leq w_{i},\text{ for all }i}X(w_{1}^{\prime},\ldots w_{r}^{\prime}).$
When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$. For any $s\in S$, the Zariski closure $\overline{BsB}=B\langle s\rangle B$ is the parabolic subgroup corresponding to the subgroup $\langle s\rangle\subset W$, cf. [DMR, Corollary 2.2.10], so the compactified Deligne–Lusztig variety $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is always smooth. To simplify the notations, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$. The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.
Furthermore, denote by $\alpha:S^{*}\to W$ be the natural projection morphism of monoids. Let $\preceq$ be the Bruhat order in $S^{*}$ and let $\ell^{S^{*}}:S^{*}\to\mathbb{N}\cup\{0\}$ be the Bruhat length.
@@ 2.2. The Lang map
We consider an alternative description of Deligne–Lusztig varieties via the Lang map. Using this description, we show that $G^{F}$ acts transitively on the set of connected components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Denote by $G^{r}$ the $r$-fold product $G\times\cdots\times G$ over ${\rm Spec}\ k$. This is again a $k$-group scheme.
@@
Let $\mathcal{L}_{1}:G^{r}\to G^{r}$ be the Lang map associated to $F_{1}$ given by $x\mapsto x^{-1}F_{1}(x)$ for $x\in G^{r}$. In other words, we have
@@
Furthermore, $G^{F}$ acts transitively on the irreducible components of $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$). In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equidimensional.
@@
By Lang’s theorem as formulated in [CabanEngue, Theorem 7.1], we deduce that $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$. The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the free $({B})^{r}$-action from the right, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action. It follows that $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is a closed subscheme of $\left(G/B\right)^{r}$. Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via projection to the first $r$-entries. The isomorphisms in this lemma then follows from verifying the definitions, cf. [DL, §1.11].
@@
$G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\overset{\sim}{\longrightarrow}\mathcal{B}$
is irreducible. Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Thus $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of irreducible components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
@@
The $G^{F}$-action on $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is not free. In $G/B$, the point ${B}$ already has nontrivial stabiliser $B^{F}=B(\mathbb{F}_{q})$. The point $(B,\ldots,B)$ in $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also stabilised by $B^{F}=B(\mathbb{F}_{q})$.
2.3. Criterion for connectedness
There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BonnafeRouquierOntheirred], Görtz [GoertzOntheconn], and many others. In this section we prove that $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BonnafeRouquierOntheirred].

### v3

@@ Abstract.
For connected reductive groups together with a Frobenius root $F$, we show that the cohomology of the structure sheaf and respectively the canonical sheaf for compactified Deligne–Lusztig varieties associated to an element in the free monoid generated by the simple reflections is isomorphic to that of a minimal length element in an $F$-conjugacy class in the Weyl group.
@@ 1.1. Background
Let $q$ be a power of a prime $p$, $k_{0}=\mathbb{F}_{q^{d}}$ for a positive integer $d$, and $k$ an algebraic closure of $\mathbb{F}_{p}$ containing $k_{0}$. Let $G_{0}$ be a connected reductive $k_{0}$-group with $q^{d}$-Frobenius $F_{0}:G_{0}\to G_{0}$, and let $G$ be the base change of $G_{0}$ to $k$ together with an endomorphism $F:G\to G$ such that $F^{d}=F_{0}\otimes_{k_{0}}{\rm Id}$. The group of fixed points $G^{F}$ is known as a finite group of Lie type. Examples of such groups include ${\rm GL}_{n}(\mathbb{F}_{q})$ and the subgroup of unitary matrices in ${\rm GL}_{n}(\mathbb{F}_{q^{2}})$.
Let $T\subset B$ be an $F$-stable maximal torus and an $F$-stable Borel subgroup of $G$. Denote by $W:=N(T)/T$ the corresponding Weyl group and let $S\subseteq W$ be the set of simple reflections. For any $w\in W$, the associated Deligne–Lusztig variety is denoted by $X(w)$. Throughout the introduction, we fix a reduced expression of $w$ and denote by $\overline{X}(w)$ the corresponding smooth compactification of $X(w)$ (Section˜2.1).
In [7], Deligne and Lusztig considered the $\ell$-adic cohomology of $X(w)$ and certain Galois coverings of $X(w)$ to study the irreducible $G^{F}$-representations over fields of characteristic zero. The de Rham and mod $p$ cohomology of $X(w)$ and $\overline{X}(w)$ are less understood than their $\ell$-adic counterpart. However, for ${\rm GL}_{2}(\mathbb{F}_{q})$ and $w$ the Coxeter element, the de Rham cohomology of $\overline{X}(w)$ is not mysterious because $\overline{X}(w)$ is isomorphic to the projective line.
@@
In [26], we described the cohomology of the structure sheaf of $\overline{X}(w)$ for any $w\in W$ in the case of ${\rm GL}_{n}(\mathbb{F}_{q})$. Our main theorem showed that the cohomology of the structure sheaf of $\overline{X}(w)$ is always isomorphic to that of $\overline{X}(c)$, where $c$ is a Coxeter element in a parabolic subgroup of $W$ with ${\text{supp}}(c)={\text{supp}}(w):=\{s\in S\ |\ s\leq w\}$. As a consequence of this theorem, we were able to describe the compactly supported integral $p$-adic cohomology of any Deligne–Lusztig variety $X(w)$ for ${\rm GL}_{n}(\mathbb{F}_{q})$.
@@ 1.2. Main results
The aim of this paper is to extend the main theorem of [26] to all pairs $(G,F)$ as in §1.1. After fixing a suitable pair $T\subset B$ of $F$-stable maximal torus and Borel subgroup, we may assume that $F(S)=S$ without loss of generality [9, §4.2]. In fact, when $G$ is semisimple, $F$ would be the Frobenius endomorphism (i.e. $d=1$) except for the cases for Suzuki and Ree groups, for which $d=2$ and additional assumptions on $q$ are required (Section˜5.2).
In our setting, we encounter the minimal length elements in the $F$-conjugacy classes of $W$ (Section˜4.2), which are special elements in $W$ with nice properties [10, Theorem 2.6], [19, Theorem 3.1]. When $G^{F}$ is ${\rm GL}_{n}(\mathbb{F}_{q})$, they are precisely the Coxeter elements in parabolic subgroups of $W$. In general, Coxeter elements of $W$ (in the sense of Remark˜2.3.2) are examples of minimal length elements in $F$-conjugacy classes.
Theorem˜5.1.1.
For every $w\in W$ there exists an $F$-conjugacy class $C$ in $W$ and a minimal length element $x_{0}\in C$ such that for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
@@
Note that the $F$-conjugacy class $C$ in the theorem may not contain $w$. However, for any given $w$, one can determine $x_{0}$ and $C$ explicitly by following our recipe given in the proof of Theorem˜5.1.1. We will also show that the cohomology groups in the theorem do not depend on the choice of reduced expressions of $w$ and $x_{0}$ and their corresponding smooth compactifications.
On the other hand, smooth compactifications of $X(w)$ given by different reduced expressions of $w$ are not necessarily isomorphic. To avoid this ambiguity, we work with compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ associated to a tuple of Weyl group elements, cf. Definition 2.1.1, [8, Definition 2.3.2]. This notion makes the construction of compactifications of Deligne–Lusztig varieties more straightforward and uniform. For example, setting $r=1$ yields the Zariski closure $X(\underline{w_{1}})=\overline{X(w_{1})}$ of $X(w_{1})$ in $G/B$. Setting $w_{1},\ldots,w_{r}\in S$ such that $w_{1}\cdots w_{r}\in W$ is a reduced expression yields the smooth compactification $X(\underline{w_{1}},\ldots,\underline{w_{r}})=\overline{X}(w_{1}\cdots w_{r})$ defined by Deligne and Lusztig.
In this generality, we give an irreducibility criterion for compactified Deligne–Lusztig varieties $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, generalising a result of Bonnafé and Rouquier [3]. We also show that they have $F_{p}$-rational singularities (Section˜2.4).
The structure of the paper is as follows. We fix our notations for Deligne–Lusztig varieties and compactified Deligne–Lusztig varieties and describe their geometric properties in Section˜2. In Section˜3, we show that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth when each $w_{i},i=1,\ldots,r,$ comes from a braid relation in the Weyl group $W$, and we establish relations between cohomology groups for the structure sheaf and the canonical sheaf on the associated compactified Deligne–Lusztig varieties.
Section˜4 contains the constructions of morphisms which are $\mathbb{P}^{1}$-fibrations between certain compactified Deligne–Lusztig varieties. They are associated with $F$-cyclic conjugations in the Weyl group. In Proposition 4.1.4, we relate the cohomology of the structure sheaf (resp. canonical sheaf) between these compactified Deligne–Lusztig varieties. Finally, we prove Theorem˜5.1.1 in Section˜5.
@@ 2. Compactified Deligne–Lusztig varieties
We begin with fixing notations related to Deligne–Lusztig varieties associated to a tuple of Weyl group elements. This construction has been treated in [7, §9] and [8, §2]. We review the definitions and properties that are relevant for our context. Moreover, by adapting classical proofs, we give criterions for compactified Deligne–Lusztig varieties to be irreducible and show that they are $F_{p}$-rational.
@@ 2.1. Definitions and basic properties
Let $G$, $F:G\to G$, $T\subset B$, $S\subset W$ be as in Section˜1.1. As above, we take a suitable pair of $F$-stable $T\subset B$ such that $F(S)=S$ [9, §4.2]. In the literature, the endomorphism $F$ is referred to as a Frobenius root or a Steinberg endomorphism. Denote by $\ell:W\to\mathbb{N}\cup\{0\}$ the Bruhat length and $\leq$ the Bruhat order on $W$.
Recall that by the Bruhat decomposition, the diagonal $G$-action on the product $G/B\times G/B$ yields a bijection between the Weyl group $W$ and $G\backslash\left(G/B\times G/B\right)$. The orbit corresponding to $w\in W$ is denoted by O(w){O(w)}. It can be regarded as a fibre bundle over the flag variety $G/B$ with fibres isomorphic to the Schubert cell $BwB/B$.
@@
Since $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is an iterated fibre bundle over $G/B$ with fibres isomorphic to Schubert varieties $\overline{Bw_{i}B}/B$, cf. [21, F.23], it is an integral, separated, normal, Cohen–Macaulay $k$-scheme locally of finite type. A proof of normality can be found in [8, Lemma 2.2.13]. The Cohen–Macaulay property behaves well under base change [2, Lemma 045T], so $O(\underline{w_{1}},\ldots,\underline{w_{r}})$ is Cohen–Macaulay follows from the classical result that Schubert varieties are Cohen–Macaulay (see the introduction of [18]).
Denote by $F:G/B\to G/B$ the map induced by the endomorphism $F:G\to G$ and $\Gamma_{F}\subseteq G/B\times G/B$ its graph. Since $F^{d}$ is a Frobenius endomorphism, we know that $F$ is surjective, finite ([9, Proposition 8.1.2]), separated ([13, Proposition 9.13]), and purely inseparable ([16, (3.7.6)]). Moreover, the induced map $F:W\to W$ is a group automorphism [9, Lemma 4.2.21].
@@
For any $w_{1},\ldots,w_{r}\in W$, the $k$-scheme $X(w_{1},\ldots,w_{r})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$. If for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth of dimension $\ell(w_{1})+\cdots+\ell(w_{r})$, cf. [8, Proposition 2.3.5, 2.3.6 (i)]. In general, compactified Deligne–Lusztig varieties have $F_{p}$-rational singularities, as we will see in Proposition 2.4.1.
@@
[…]
When $s_{i}\in S$ for all $i=1,\ldots,r$, the tuple $(s_{1},\ldots,s_{r})$ can be regarded as an element in the free monoid generated by $S$, which we denote by $S^{*}$. For any $s\in S$, the Zariski closure $\overline{BsB}=B\langle s\rangle B$ is the parabolic subgroup corresponding to the subgroup $\langle s\rangle\subset W$, cf. [8, Corollary 2.2.10], so the compactified Deligne–Lusztig variety $X(\underline{s_{1}},\ldots,\underline{s_{r}})$ is always smooth. To simplify the notations, given any $w=s_{1}\cdots s_{m}\in S^{*}$ with $s_{i}\in S,i=1,\ldots m$, we denote $X(s_{1},\ldots,s_{m})$ by $X^{S^{*}}(w)$ and $X(\underline{s_{1}},\ldots,\underline{s_{m}})$ by $\overline{X}^{S^{*}}(w)$. The notations $O^{S^{*}}(w)$ and $\overline{O}^{S^{*}}(w)$ for the orbits are defined analogously.
If $s_{1}\cdots s_{r}\in S^{*}$ is a reduced expression of an element in $w\in W$, then $\overline{X}^{S^{*}}(s_{1}\cdots s_{r})$ recovers the smooth compactification $\overline{X}(w)$ as defined in [7, §9.10]. Furthermore, denote by $\alpha:S^{*}\to W$ be the natural projection morphism of monoids. Let $\preceq$ be the Bruhat order in $S^{*}$ and let $\ell^{S^{*}}:S^{*}\to\mathbb{N}\cup\{0\}$ be the Bruhat length.
@@ 2.2. The Lang map
We consider an alternative description of Deligne–Lusztig varieties via the Lang map. Using this description, we show that $G^{F}$ acts transitively on the set of irreducible components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Denote by $G^{r}$ the $r$-fold product $G\times\cdots\times G$ over ${\rm Spec}\ k$. This is again a $k$-group scheme.
@@
The Lang map $\mathcal{L}_{1}$ associated to $F_{1}$ is the endomorphism $x\mapsto x^{-1}F_{1}(x)$ of $G^{r}$. In other words, we have
@@
Moreover, $G^{F}$ acts transitively on the irreducible components of $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$). In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equidimensional.
@@
By Lang’s theorem as formulated in [5, Theorem 7.1], we deduce that $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$. The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the free $({B})^{r}$-action from the right, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action. It follows that $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is a closed subscheme of $\left(G/B\right)^{r}$. Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via projection to the first $r$-entries. The isomorphisms in this lemma then follows from verifying the definitions, cf. [7, §1.11].
@@
$G^{F}\backslash\mathcal{L}_{1}^{-1}(\mathcal{B})\overset{\sim}{\longrightarrow}\mathcal{B}$
is irreducible. Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash\mathcal{L}_{1}^{-1}(\mathcal{B})\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Thus $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of irreducible components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
@@
The $G^{F}$-action on $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is not free. In $G/B$, the point B{B} already has nontrivial stabiliser $B^{F}=B(\mathbb{F}_{q})$. The point $(B,\ldots,B)$ in $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also stabilised by $B^{F}=B(\mathbb{F}_{q})$.
2.3. Criterion for irreducibility
There is a simple criterion for the irreducibility of classical Deligne–Lusztig varieties, proved by Lusztig, Bonnafé–Rouquier [3], Görtz [15], and many others. In this section we show that $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is irreducible when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [3].